# Privacy-Aware Semantic Cache for Large Language Models

### v2

@@ Abstract.
Large Language Models (LLMs) like ChatGPT and Llama2 have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters where inference demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries which constitute about 31% of the total queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries, leading to unacceptable false hit-and-miss rates.
This paper introduces MeanCache, a user-centric semantic cache for LLMs that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. Existing caching solutions for LLMs raise privacy and scalability concerns and perform wasteful query requests. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model across LLM users without violating privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower false hit rates. MeanCache compresses the embedding dimensions to minimize cache storage and also finds the optimal cosine similarity threshold. Our experiments benchmarked against the state-of-the-art caching method, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions. It also reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%.
Federated Learning, Cache, Hit, Miss, Rate, ChatGPT, Llama, Semantic, Search, Large Language Models, LLMs, GPT, Machine Learning, Privacy, Embeddings, Cosine Similarity, Client
@@ 1. Introduction
Large Language Models (LLMs) like ChatGPT (cha, [n.d.]), Google Bard (bar, [n.d.]), Claude (cla, [n.d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in applications ranging from search engines to conversational agents. LLMs are increasingly integrated into platforms like the Perplexity AI search engine  (per, [n.d.]b), Rabbit OS (rab, [n.d.]), Arc browser (arc, [n.d.]), and Bing Chat (bin, [n.d.]). For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n.d.]).
@@
Consequently, LLM-based services charge users and limit query rates (per, [n.d.]a; ope, [n.d.]). Prior studies report that users frequently submit similar queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002)—approximately 33% of search engine queries being resubmitted (Markatos, 2001). We investigate if a similar pattern exists with queries to LLMs. Our analysis of a large corpus of real-world user conversations with ChatGPT reveals that approximately 31% of such queries have a high similarity with a previously submitted query 2. Thus, there is a potential to reduce the query inference cost of LLMs by up to one-third.
@@
Existing semantic caches have several limitations. First, they demand a significantly large central cache to store the queries and responses of all users, which is unscalable and violates users’ privacy. Second, they incur the network cost of sending a user query to the server even if there is a cache hit. An end user will still be charged for the query even if the query is served from the server cache. Third, they use a single server-side embedding model to find the semantic similarity among queries, which does not generalize to each user’s querying patterns For instance, Google Keyboard (gbo, [n.d.]) adapts to each user’s unique writing style and embeds such personalized behaviors to enhance the accuracy of the next word prediction model. Fourth, they employ Llama-2 to enhance the accuracy of semantic matching (Bang, 2023); however, in practice, such models perform billions of operations to generate embeddings, offsetting the benefits of the cache.
@@
To find a semantic match between a new query and cached queries, prior work (Bang, 2023) advocates for a larger model, such as Llama 2, to generate highly accurate embeddings but requires billions of operations. Such models are not suitable for resource-limited users’ devices. Previous work has shown that a smaller model can achieve performance comparable to larger models on custom tasks  (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). MeanCache uses smaller embedding models such as MPNet (Song et al., 2020) to generate embeddings for semantic matching locally.
@@
MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching, while still outperforming the state-of-the-art GPTCache.
@@ 3.2. Transformer Architecture
The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer
helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
@@ 4. MeanCache’s Design
MeanCache is a user-centric semantic caching system optimized for user-side operation. Figure 2 illustrates the workflow of MeanCache for similar queries. When a user submits a query to an LLM web service with MeanCache enabled, MeanCache computes the query embeddings. These embeddings are then matched with the embedding of the cached queries using cosine similarity. If MeanCache locates a similar query within the cache, the response is retrieved from the local cache and returned to the user. Otherwise, MeanCache forwards the query to the LLM web service to obtain the response. The query and its response and embeddings are then stored in the cache.
The key contribution of MeanCache comes from harnessing the collective intelligence of multiple users to training a semantic similarity model and its user-centric design that addresses both privacy and scalability issues. To achieve these, MeanCache takes the following design decisions. It employs a small embedding model with lower computational overhead than large Language Learning Model (LLM) based embedding models, such as Llama 2. It uses federated learning (FL) for collaborative training to fine-tune the embedding model without ever storing user data on a central server. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. To reduce storage and memory overhead and expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). MeanCache leverages user feedback (i.e., forcing LLM request instead of the cache query) to adapt the threshold for cosine similarity. It then shares the updated threshold via FL to determine the optimal threshold across multiple users. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache on clients.
@@ 4.1. FL Based Embedding Model Training
After finding the global optimal threshold and the global embedding model, the server then redistributes the updated embedding model to the users for the next round of FL training. This process is repeated over several to improve the semantic matching accuracy of the MeanCache (i.e., lower false hits and false misses). This approach ensures that the user’s privacy is maintained (i.e., clients queries are never stored at the central server), and the collective intelligence of multiple users is leveraged to improve the performance of the MeanCache. After the completion of the FL training, each client will have access to a fine-tuned embedding model ($W_{global}^{t+1}$) to generate high-quality embeddings which can capture the complex semantics of a user query. For example, MeanCache’s embedding model can discern the semantic similarity between ”Draw a line plot in Python” and ”Plot a linear graph using Python”, and serve such queries locally, as depicted in Figure 2.
@@ 4.2. Cache Population and MeanCache Implementation
Once the embedding model is trained within MeanCache on each user, it is deployed as depicted in Figure 2. Initially, when a new user starts using MeanCache, the local cache is vacant. During these interactions, if a user query’s response is not found in the cache, the request is forwarded to the LLM web service to retrieve the response, which is then inserted in the cache. If MeanCache finds a semantically similar query for any of the following queries, it responds to the user’s similar queries by retrieving the response from the local cache.
@@ 5. Evaluation
Our evaluation answer the following research questions.
  • Is it possible to train an embedding model in a privacy-preserving manner without the need for centralized data?
  • Besides preserving user query privacy, how does MeanCache perform in comparison to baseline in terms of performance metrics?
  • What effect does the cosine similarity threshold have on the performance of MeanCache?
  • What level of overhead does MeanCache impose on the user side? Is it possible to reduce the embedding dimension to save storage space while still outperforming the baseline?
  • GPTCache (Bang, 2023) suggests using Llama 2 to generate embeddings to improve semantic matching. Is it feasible to use Llama 2 to compute embeddings at the user side for semantic matching?
@@ 5.1. Evaluation Settings
We evaluate MeanCache using the GPTCache dataset. The dataset is partitioned into training, testing, and validation subsets. The training and validation datasets are randomly distributed among the clients, with each client receiving non-overlapping data points. During local training, each client utilizes its training dataset to update its local embedding model and employs the validation dataset to determine the optimal threshold for cosine similarity (Section 5.3). The testing dataset, located at the central server, facilitates a fair comparison between MeanCache and GPTCache (Bang, 2023).
@@ 5.1.3. Evaluation Metrics
In caching systems, the efficacy has traditionally been gauged by cache hit-and-miss rates. A cache hit implies the data or query is retrieved from the cache, whereas a cache miss indicates the opposite. Semantic caching introduces a nuanced classification: true and false hits, alongside true and false misses. A true hit signifies a correct match between a query and a similar cached query, whereas a false hit is an incorrect match with a non-similar cached query.
Thus, semantic caching complexity renders traditional hit/miss metrics potentially misleading. For example, a query might incorrectly match with an irrelevant cached query (deemed a hit traditionally) due to semantic matching. This necessitates more sophisticated metrics to accurately assess a semantic cache’s performance. We adopt precision, recall, F1 score, and accuracy, for a comprehensive evaluation of MeanCache against GPTCache. These metrics are defined as follows:
@@
Figure 5. FL training enhances MPNet’s ability to generate high-quality embeddings by leveraging users’ private data, thereby improving MeanCache’s query matching precision from 0.74 to 0.85, demonstrating its effectiveness.
@@
Figure 6. FL training boosts MeanCache’s user query matching precision with Albert embedding model from 0.74 to 0.81, ensuring privacy.
@@ 5.2. Privacy Preserving Embeddings Model Training
Figure 7. Extremes in threshold harm semantic matching; MeanCache autonomously optimizes MPNet’s threshold at 0.83 for an F score of 0.89.
@@ 5.3. Cosine Similarity Threshold Impact on Semantic Matching
Figure 11. Confusion matrices for MeanCache and GPTCache on 1000 queries comprising 700 unique and 300 duplicate queries. Among the 700 unique queries, MeanCache produces only 89 false hits, while GPTCache generates a significantly larger number of false hits, 233 in total. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
Metrics GPTCache MeanCache MeanCache
@@
F score 0.56 0.73 0.68
Precision 0.52 0.72 0.66
Recall 0.85 0.78 0.77
Accuracy 0.72 0.85 0.81
Table 1. MeanCache outperforms GPTCache (Baseline)
@@
5.4. MeanCache Comparison with Baseline (GPTCache)

### v3

# MeanCache: User-Centric Semantic Cache for Large Language Model Based Web Services

@@ Abstract.
Large Language Models (LLMs) like ChatGPT and Llama have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters, where inference demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries, which constitute about 31% of the total queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries nor do they operate on contextual queries, leading to unacceptable false hit-and-miss rates.
This paper introduces MeanCache, a user-centric semantic cache for LLM-based services that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model without violating user privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower false hit rates. MeanCache also encodes context chains for every cached query, offering a simple yet highly effective mechanism to discern contextual query responses from standalone. Our experiments benchmarked against the state-of-the-art caching method, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions while performing even better on contextual queries. It also reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%.
Federated Learning, Caching, Large Language Models, Cache, Hit, Miss, Rate, ChatGPT, Llama, Semantic, Search, Large Language Models, LLMs, GPT, Machine Learning, Privacy, Web, Embeddings, Cosine Similarity, Client, Prompt, Query,
@@ 1. Introduction
Large Language Models (LLMs) like ChatGPT (cha, [n. d.]), Google Bard (bar, [n. d.]), Claude (cla, [n. d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in applications ranging from search engines to conversational agents. LLMs are increasingly integrated into platforms like the Perplexity AI search engine, Rabbit OS (rab, [n. d.]), Arc browser (arc, [n. d.]), and Bing Chat (bin, [n. d.]). For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n. d.]).
@@
Consequently, LLM-based services charge users and limit query rates (per, [n. d.]; ope, [n. d.]). Prior studies report that users frequently submit similar queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002)—approximately 33% of search engine queries being resubmitted (Markatos, 2001). We investigate if a similar pattern exists with queries to LLMs. Our analysis of a large corpus of real-world user conversations with ChatGPT reveals that approximately 31% of such queries have a high similarity with a previously submitted query, as shown in Section 2. Thus, there is a potential to reduce the query inference cost of LLMs by up to one-third.
@@
Existing semantic caches have several limitations. First, they demand a significantly large central cache to store the queries and responses of all users, which is unscalable and violates users’ privacy. Second, they incur the network cost of sending a user query to the server even if there is a cache hit. An end user will still be charged for the query even if the query is served from the server cache. Third, they use a single server-side embedding model to find the semantic similarity among queries, which does not generalize to each user’s querying patterns. For instance, Google Keyboard (gbo, [n. d.]) adapts to each user’s unique writing style and embeds such personalized behaviors to enhance the accuracy of the next word prediction model. Fourth, they employ Llama-2 to enhance the accuracy of semantic matching (Bang, 2023); however, in practice, such models perform billions of operations to generate embeddings, offsetting the benefits of the cache. Lastly, they are only effective on standalone LLM queries, resulting in unbearably high false hit rates for contextually different but semantically similar queries.
@@
To find a semantic match between a new query and cached queries, MeanCache uses smaller embedding models such as MPNet (Song et al., 2020) to generate embeddings for semantic matching locally. Previous work has shown that a smaller model can achieve performance comparable to larger models on custom tasks (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). Prior work (Bang, 2023) proposes using large language models, such as Llama 2, to generate embeddings. However, these models require billions of operations, making them unsuitable for embeddings generation.
Due to different contexts around queries, LLM may return different responses for semantically similar queries. For such queries, MeanCache also records the contextual chain, parent queries already in the cache, for a given query. To find a response for a contextual query, MeanCache verifies the context of a contextual query by matching a given query’s context with the cached query’s context chain to accurately retrieve responses for previously seen contextually different but semantically similar queries from the cache.
@@
MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F-score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s performance on contextual queries is even more impressive when compared to GPTCache (baseline). For contextual queries, MeanCache achieves a 25% higher F-score and accuracy, and a 32% higher precision over the baseline. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching while still outperforming the state-of-the-art GPTCache.
Artifact Availability: MeanCache is implemented in the Flower FL framework (Beutel et al., 2020). The complete source code and contextual queries dataset will be available upon acceptance.
@@ 3.2. Transformer Architecture
The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
@@ 4. MeanCache’s Design
MeanCache is a user-centric semantic caching system optimized for user-side operation. Figure 2 illustrates the workflow of MeanCache for similar queries. Algorithm 1 further explains MeanCache’s querying and population process programmatically. When a user submits a query to an LLM web service with MeanCache enabled, MeanCache computes the query embeddings (Line  3). These embeddings are then matched with the embedding of the cached queries using cosine similarity (Line  4). For every similar query found within the cache, MeanCache analyzes the context chain for every query (Line  7) and matches its embedding with the conversational history (Line  9) of the submitted query. If MeanCache finds a similar query with a similar context chain, the response is retrieved from the local cache and returned to the user (Line  13). Otherwise, MeanCache forwards the query to the LLM web service to obtain the response. The query and its response and embeddings are then stored in the cache (Line  18).
The key contribution of MeanCache comes from harnessing the collective intelligence of multiple users to training a semantic similarity model and its user-centric design that addresses both privacy and scalability issues. To achieve these, MeanCache takes the following design decisions. It employs a small embedding model with lower computational overhead than large Language Learning Model (LLM) based embedding models, such as Llama 2. It uses federated learning (FL) for collaborative training to fine-tune the embedding model without ever storing user data on a central server. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. To handle contextual queries, MeanCache includes contextual chain information in its cache against every query to identify if the cached response for a query is only applicable under a specific context. This design is capable of handling long contextual chains. To reduce storage and memory overhead and expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). MeanCache leverages user feedback (i.e., forcing LLM request instead of the cache query) to adapt the threshold for cosine similarity. It then shares the updated threshold via FL to determine the optimal threshold across multiple users. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache on clients.
Algorithm 1 MeanCache Workflow for LLM Queries
1:Input: User query $Q$, User Query Context $C_{q}$
2:Output: Response $R$
3:Compute query embedding $E_{Q}$
4:$similar\_queries\leftarrow$ FindSimilarQueriesinCache($E_{Q}$)
5:$context\_match\leftarrow$ False
6:for each query $C_{i}$ in $similar\_queries$ do
7:    $ctx\leftarrow$ the entire context chain of $C_{i}$
8:    $E_{ctx}\leftarrow$ Retrieve embedding of $ctx$
9:    $E_{cq}\leftarrow$ Retrieve embedding of $C_{q}$
10:    Compute cosine similarity $s_{i}$ between $E_{cq}$ and $E_{ctx}$
11:    if $s_{i}\geq$ context similarity threshold then
12:        $context\_match\leftarrow$ True
13:        Retrieve response $R$ from cache for $C_{i}$
14:        return $R$
15:    end if
16:end for
17:if not $context\_match$ then
18:    $R\leftarrow$ LLMResponseAndEnrollInCache($Q$, $E_{Q}$ , $C_{q}$)
19:end if
20:return $R$
@@ 4.1. FL Based Embedding Model Training
After finding the global optimal threshold and the global embedding model, the server then redistributes the updated embedding model to the users for the next round of FL training. This process is repeated over several to improve the semantic matching accuracy of the MeanCache (i.e., lower false hits and false misses). After the completion of the FL training, each client will have access to a fine-tuned embedding model ($W_{global}^{t+1}$) to generate high-quality embeddings that can capture the complex semantics of a user query.
@@ 4.2. Cache Population and MeanCache Implementation
Once the embedding model is trained within MeanCache on each user, it is deployed as depicted in Figure 2. Initially, when a new user starts using MeanCache, the local cache is vacant. During these interactions, if a user query’s response is not found in the cache, the request is forwarded to the LLM web service to retrieve the response, which is then inserted in the cache. If MeanCache finds semantically similar queries in the cache for any of the following queries from the user, it analyzes the context chain for every similar cached query and matches its embedding with the conversational history of the submitted query. If MeanCache finds a similar query with a similar context chain, the response is retrieved from the local cache and returned to the user. Otherwise, MeanCache forwards the query to the LLM web service to obtain the response. The query and its response and embeddings are then stored in the cache.
@@ 5. Evaluation
Our evaluation answers the following research questions.
  • Is it possible to train an embedding model in a privacy-preserving manner without the centralized data (section 5.2)?
  • What effect does the cosine similarity threshold have on the performance of MeanCache (section 5.3)?
  • Besides preserving user query privacy, how does MeanCache perform in comparison to baseline in terms of performance metrics (section 5.4)?
  • How accurately does MeanCache retrieve contextual queries from the cache (section 5.5)?
  • What level of overhead does MeanCache impose on the user side? Is it possible to reduce the embedding dimension to save storage space while outperforming the baseline (section 5.6)?
  • GPTCache (Bang, 2023) suggests using Llama 2 to generate embeddings to improve semantic matching. Is it feasible to use Llama 2 to compute embeddings at the user side for semantic matching (section 5.7)?
@@ 5.1. Evaluation Settings
We evaluate MeanCache using the GPTCache dataset. The dataset is partitioned into training, testing, and validation subsets. The training and validation datasets are randomly distributed among the clients, with each client receiving non-overlapping data points. During local training, each client utilizes its training dataset to update its local embedding model and employs the validation dataset to determine the optimal threshold for cosine similarity (Section 5.3). The testing dataset, located at the central server, facilitates a fair comparison between MeanCache and GPTCache (Bang, 2023). Since there does not exist any dataset of contextual queries, we generate a synthetic dataset using GPT-4 consisting of 450 queries, including duplicates, non-duplicates, and contextual queries, to evaluate MeanCache performance on contextual queries.
@@ 5.1.3. Evaluation Metrics
In caching systems, the efficacy has traditionally been gauged by cache hit-and-miss rates. A cache hit implies the data or query is retrieved from the cache, whereas a cache miss indicates the opposite. Semantic caching introduces a nuanced classification: true and false hits, alongside true and false misses. A true hit signifies a correct match between a query and a similar cached query, whereas a false hit is an incorrect match with a non-similar cached query. A true miss signifies when a query does not have a similar cached query, whereas a false miss is when a query has a similar cached query but is not received from the cache. Thus, traditional hit/miss metrics are potentially misleading in semantic caches. For example, a query might incorrectly match with an irrelevant cached query (deemed a hit traditionally) due to semantic matching. We adopt precision, recall, F score, and accuracy for a comprehensive evaluation of MeanCache against GPTCache (baseline). These metrics are defined as follows:
@@
Figure 5. In MeanCache, MPNet’s FL training helps generate high-quality embeddings by leveraging users’ private data, thereby improving query matching precision.
@@
Figure 6. FL training boosts MeanCache’s user query matching precision with Albert embedding model.
@@ 5.2. Privacy Preserving Embeddings Model Training
Figure 7. MeanCache automatically optimizes MPNet’s threshold.
@@ 5.3. Cosine Similarity Threshold Impact on Semantic Matching
Figure 11. Confusion matrices for MeanCache and GPTCache on 1000 queries comprising 700 unique and 300 duplicate queries. Among the 700 unique queries, MeanCache produces only 89 false hits, while GPTCache generates a significantly larger number of false hits, 233 in total.
Metrics Standalone Queries Contextual Queries
GPT- Mean- Mean- GPT- MeanCache
Cache Cache Cache Cache
@@
F score 0.56 0.73 0.68 0.67 0.93
Precision 0.52 0.72 0.66 0.66 0.98
Recall 0.85 0.78 0.77 0.71 0.79
Accuracy 0.72 0.85 0.81 0.61 0.86
Table 1. MeanCache outperforms GPTCache (baseline) on both standalone and contextual queries.
@@
5.4. MeanCache Comparison with Baseline
@@
5.5. Contextual Queries
(a) Ideally, all 100 queries should result in misses. However, GPTCache incorrectly produces 54 false hits, while MeanCache yields only 3.
(b) MeanCache yields 8% more true hits than the baseline.
Figure 13. Performance on Contextual Queries: MeanCache vs. Baseline. (a) reports MeanCache’s fewer false hits 3 vs. 54 of GPTCache. (b) reports higher true hits by MeanCache.
(a) MeanCache
(b) GPTCache (baseline)
Figure 14. MeanCache only reports three false hits compared to 54 false hits by GPTCache. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
The end-user interaction with an LLM service mainly involves two types of user queries: standalone (e.g., Q1 Draw a line in Python?) and follow-up or contextual queries (e.g., Q2 Change the color to red). While standalone queries can be processed independently, contextual queries require prior context for accurate responses. Suppose that both Q1 and Q2 with their responses are cached. The user sends a new query, Q3 Draw a circle?, followed by Q4 Change the color to red. Q4 is semantically similar to Q2 in the cache but refers to a different context. GPTCache incorrectly identifies this as a cache hit, leading to an inaccurate response. MeanCache addresses this issue by verifying the context chain (Algorithm 1) of semantically matched queries. This approach ensures that contextual queries (Q4) result in a true cache miss, prompting a request to the LLM service for an appropriate response.
On the dataset of 450 contextual queries (see Section 5.1.1), first, we populate the cache (MeanCache and baseline) with 200 queries (100 standalone and 100 contextual queries of the standalone queries). Next, we send 150 duplicate queries (75 standalone + 75 contextual) and 100 non-duplicate queries (a total of 250 queries) to the cache-enabled LLM. Figure 13 shows the true label (whether the query should be returned from the cache or not) and the corresponding GPTCache and MeanCache performance (predicted label). Note that in Figure 13(a), all the queries should be answered by the LLM; in other words, there should be no hits. However, GPTCache has 54 false hits, while MeanCache has only three false hits. This is also shown in the confusion matrix in Figure 14. Table 1 (Column-3) summarizes the comparative results. MeanCache outperforms GPTCache by over 25% in both F-Score and accuracy. Additionally, MeanCache achieves 32% higher precision compared to the baseline.
Summary. GPTCache’s low performance stems from its inability to consider contextual information, leading to high false hit rates. In contrast, MeanCache demonstrates significantly superior handling of contextual queries, with a 25% improvement in accuracy over the baseline.